# User-Driven Value Alignment: Understanding Users’ Perceptions and Strategies for Addressing Biased and Discriminatory Statements in AI Companions

### v1

@@ 6. DISCUSSION
We have taken initial steps to explore an emerging phenomenon in which users of LLM-based AI companions actively identify, challenge, and attempt to correct AI outputs they perceive as harmful, aiming to guide the AI to better align with their values. Below, we summarize the unique contributions of user-driven value alignment, raise open questions on how to support it, and discuss its challenges and limitations.
@@ 6.1. What is User-Driven Value Alignment?
First, in user-driven value alignment, users have greater agency over the AI alignment process. Existing work on value alignment is primarily expert-driven, where technical experts design guardrails in a top-down manner to prevent harmful AI outputs (Askell et al., 2021; Bai et al., 2022; Sorensen et al., 2024; Guo et al., 2024). Even when humans are involved (e.g., via approaches like Reinforcement Learning with Human Feedback (Christiano et al., 2017a)), they are often involved in a manner that lacks of meaningful agency. Our research shows that when re-aligning biased AI companions, users took the lead in identifying harmful machine behaviors, deciding which behaviors to address, and choosing how to address them. They employ various strategies—including technical adjustments, argumentative discussions, and character modifications—to mitigate AI bias. In this way, users transition from being passive consumers to active participants, directly participating in shaping AI behavior.
@@
Second, user-driven value alignment is grounded in real-world contexts. While existing expert-driven approaches provide valuable insights, they often fail to fully anticipate and address the nuanced and diverse needs that users encounter in real-world contexts (Mirowski et al., 2024). These approaches may lack the flexibility needed to adapt to the complexities of human-AI interactions, resulting in gaps in effectively aligning AI behavior with the varied expectations of everyday users (Klingefjord et al., 2024). In contrast, our study reveals that when users try to re-align biased AI companions, they can identify specific statements they deem to be harmful and respond to those statements based on the unique contexts in which they are situated, making the alignment process more relevant.
@@
Third, user-driven value alignment centers around personal and community expectations, values, and norms. Previous literature has highlighted how existing top-down, one-size-fits-all value alignment processes—often termed global cultural value alignment—may overlook or even conflict with individual expectations and local community norms (Mirowski et al., 2024; Varshney, 2023). Our research echoes these concerns, emphasizing the importance of tailoring the alignment process to better reflect the unique values and needs of users and their communities. Indeed, when some users detected biased statements, they deliberately chose not to correct the AI through technical means (e.g., directly modifying the AI settings) because doing so would disrupt their experiences and expectations with the AI companions (Section 5.4.1).
@@
Finally, user-driven value alignment supports iterative human-AI interactions. Unlike traditional user-driven algorithm auditing, user-driven value alignment involves users not only identifying AI biases but also actively participating in correcting and mitigating these harms. In the past, users struggled to directly intervene and modify system behavior (Shen et al., 2021a; Li et al., 2023); however, our study shows that users can potentially influence the memory and future behavior of a specific AI companion through iterative interactions. This ease of engagement fosters the development of user-driven value alignment, where users play a crucial role in shaping AI systems, and highlights the potential for building AI systems that better align with user values through ongoing processes.
6.2. Supporting User-Driven Value Alignment
@@
We discuss design implications for supporting more meaningful, effective, and safer user-driven value alignment.
@@
Our research reveals that when users encounter biases and discrimination in AI companions, they often share the incidents via social media platforms. Some more experienced users also actively share strategies for addressing those harms and re-aligning their companions. Our research starts by collecting complaining posts users shared on a wide range of social media platforms. Through these activities, users—particularly those from marginalized social groups—engage in community-building efforts similar to what Nancy Fraser describes as “counterpublics” (Fraser, 2014; Shen et al., 2021a). In those spaces, members of often marginalized social groups collectively participate in their own form of sensemaking, opinion formation, and consensus building. However, there is currently a lack of dedicated user communities focused on addressing biases in AI companions, which results in most users not having access to the strategies shared by more experienced users. Therefore, tools that support user-driven value alignment should consider how to integrate community-driven value alignment to more effectively assist users when they encounter harmful behavior from AI companions.
@@ 6.2.2. Enhancing effective user-driven value alignment via collaboration between experts and users
A combined approach that integrates expert-driven and user-driven value alignment should be considered. Expert-driven alignment provides a solid foundational framework, while user-driven alignment allows ongoing, context-sensitive adjustments. By merging these two approaches, AI systems can achieve a more dynamic and adaptable value alignment process. Empowering users through value-sensitive learning is key—AI should continuously learn and iterate based on user feedback, regularly updating personalized knowledge bases to better reflect diverse user values. When users flag harmful behaviors or provide improvement suggestions, this information should be integrated into the model’s optimization process. Future designs should prioritize user experiences and feedback, particularly from marginalized groups, to help AI more comprehensively understand different values, thereby reducing the likelihood of generating biased responses.
@@ 6.2.3. Facilitating safe user-driven value alignment via platform policies and affordances
First, for highly anthropomorphic AI agents like AI companions, platforms should have stronger constraints on their behavior to regulate what they should or should not express. This might be achieved by improving the instruction-following capabilities of the underlying LLM (Guo et al., 2024; Mu et al., 2023), optimizing output filtering mechanisms (Glukhov et al., 2023), or establishing a more comprehensive user-specific interaction database to store and reference positive and negative interaction examples.
Second, platforms should be designed to ensure that users are fully aware of the resources available in case of a harmful value misalignment. It’s crucial that users know they have the option to re-align their AI companions. This approach goes beyond simply providing re-alignment tools; it emphasizes the importance of clearly communicating these options to users, empowering them to take corrective action rather than leaving the platform out of frustration.
Third, platforms should offer robust support when users need to re-align their AI companions, rather than leaving them to rely solely on intuition—especially since some strategies may be less effective, particularly for novice users. This support could include value alignment assistants, platforms for sharing effective strategies, and other resources. These tools should be tailored to the user’s conceptual understanding of the AI companion. For example, discussing technical details like training data might not resonate with users who view their AI companion as a “baby” and could disrupt their experience, as mentioned in Section 5.3.2. Additionally, it’s essential that platforms clearly communicate the steps, tools, and resources available for re-alignment, ensuring that users understand how to restore their relationship with the AI after a misalignment rather than abandoning the application entirely.
6.3. Limitations and Challenges of User-driven Value Alignment
This section will discuss the concerns and limitations related to the emerging phenomenon of user-driven value alignment and raise some open questions.
@@ 6.3.1. Concerns about malicious users
Malicious users may pose risks to AI systems, a concern that has been a focus in the field of responsible AI. A typical example is Microsoft’s chatbot Tay, which quickly exhibited racist behavior after interacting with users on Twitter (Schwartz, 2019). While our research underscores the many advantages of enabling user-driven value alignment, it is crucial to acknowledge the associated risks and challenges. As discussed in Sections 5.4, some users may maliciously exploit the AI’s Chameleon Effect (Chartrand and Bargh, 1999), where the AI mimics human behavior to manipulate and contaminate the training data. They might use insulting or threatening language, causing the AI to replicate similar harmful behavior in future interactions. Indeed, one key advantage of expert-driven, top-down value alignment is its ability to prevent harm through strict oversight, with platforms being held accountable for ensuring that AI systems operate safely and ethically. This approach centralizes responsibility, allowing for more consistent and controlled implementation of safeguards. Shifting this responsibility to users and smaller user communities can inevitably introduce potential risks. To mitigate these risks, it is essential to explore mechanisms that combine both expert-driven and user-driven value alignment (Section 6.2.2) and to foster greater community accountability (Section 6.2.1).
@@
6.3.2. Concerns about potential harms and emotional burdens on users
When discussing the direct involvement of users in aligning harmful AI systems, it’s important to consider the potential negative impacts, such as psychological harm and emotional burdens, that this process might impose on users. Existing literature on content moderation has documented the harmful effects and emotional labor that content moderators endure when dealing with harmful social media content (Dosono and Semaan, 2019; Steiger et al., 2021; Wohn, 2019). Our research highlights that with the development of LLMs when users interact with highly anthropomorphic AI agents, such as AI companions, they often develop parasocial relationships, which could potentially intensify the harms noted in the existing literature (Section 5.1). Therefore, existing harm reduction tools and systems (Karunakaran and Ramakrishan, 2019; Steiger et al., 2022) need to be adaptively adjusted and strengthened to better support users’ needs.
6.3.3. When the design implication is not to re-align?
@@
There are questions about whether we should choose to re-align certain harmful AI systems at all. User-driven value alignment can be considered a type of “repair” work (Velkova and Kaun, 2021) that users perform in their daily interactions with AI companions to reduce biases. However, researchers have also pointed out that in some cases, “refusal”—rather than “repair”—should be considered a more appropriate response (Ganesh and Moss, 2022; Zong and Matias, 2024). In our study, although users enjoy interactions with AI companions overall, as the frequency of chats increases, users may become overly reliant on AI companions for emotional support, neglecting real-life relationships (Zimmerman, 2009). This is especially concerning, considering some users are already vulnerable. We must be cautious of AI companions becoming accomplices in reinforcing social inequalities, further isolating marginalized groups from real-life interactions. This raises the question: At what point should we consider imposing strict policies to regulate AI systems (especially those highly anthropomorphic AI agents)? When the design implication is not to re-align (Baumer and Silberman, 2011)?
@@ 6.4. Limitations and Future Work
As the first work conceptualizing and studying user-driven value alignment, this research is highly exploratory in nature, and there are a few important limitations. First, our research relies on self-reported data from interviews and social media posts, which biases may influence. The study primarily focuses on users’ perspectives, so insights from other stakeholders involved in the value alignment process would be valuable for gaining a more holistic understanding of the phenomenon. Second, the qualitative nature of the study and its limited sample size limit our ability to generalize the findings broadly. Future research should include follow-up studies using different methods to better understand users’ interactions with AI companions and the effectiveness of various alignment strategies. Third, this study focuses on AI companion, a specific type of AI application that may not represent all AI systems. It focuses on biases and discriminatory AI behaviors, which may not encompass all the problematic behaviors that a value alignment process should address. It will be valuable to explore whether and how useful the concept of user-driven value alignment is in terms of studying other types of value alignments in different AI systems. Fourth, this study emphasizes biases and harmful behaviors in LLMs; future research should consider multimodal risks and improve mechanisms for detecting and correcting multimodal biases.
@@ Appendix A Information of participants in the study
Appendix B Higher-level themes
We summarize the higher-level themes we identified through a reflexive thematic analysis approach in Table 3. Due to the limitation of word counts and space, we do not include the 41 first-level themes and 675 codes in the table.
| Third-level themes | Second-level themes |
| Types of discriminatory statements | Misogyny |
| as perceived by users | LGBTQ+ bias |
| Appearance bias | |
| Ableism | |
| Racism | |
| Socioeconomic bias | |
| Users’ conceptualization of AI companion behavior | AI as Machine |
| AI as Baby | |
| AI as Cosplayer | |
| User-driven alignment strategies for addressing | Technical Strategies |
| discriminatory statements in AI companions | Argumentative strategies |
| Character Strategies |

### v2

@@ 6. DISCUSSION
We have taken steps to explore an emerging phenomenon in which users of LLM-based AI companions actively identify, challenge, and attempt to correct AI outputs they perceive as harmful, aiming to guide the AI to align with their values. Below, we summarize the unique contributions of user-driven value alignment and discuss its challenges and limitations.
@@ 6.1. What is User-Driven Value Alignment?
First, in user-driven value alignment, users have greater agency over the AI alignment process. Existing work on value alignment is primarily expert-driven, where technical experts design guardrails to prevent harmful outputs in LLM-based systems (Askell et al., 2021; Bai et al., 2022; Sorensen et al., 2024; Guo et al., 2024; Sun et al., 2023; Meadows et al., 2024). Even when non-experts are involved (e.g., through bottom-up approaches like reinforcement learning (Christiano et al., 2017; Pan et al., 2023) or interactive goal elicitation algorithms (Mechergui and Sreedharan, 2024)), their role is often limited to simply providing training data or feedback, without opportunities for controlling the alignment process. In contrast, user agency is a crucial value in HCI (Bennett et al., 2023; Coyle et al., 2012; Madary, 2022; Feng et al., 2024; Lukoff et al., 2021). Our research shows that when re-aligning biased AI companions, users take the lead in identifying harmful machine behaviors, deciding which behaviors to address, and choosing how to address them. They employ various strategies—including technical, argumentative, and character strategies—to mitigate AI bias. In this way, users transition from being passive consumers to active participants, directly participating in shaping AI behavior.
@@
Second, user-driven value alignment is grounded in real-world contexts. While existing approaches offer valuable insights, they often fall short in anticipating and addressing the complex, diverse needs and challenges that users face in real-world contexts (Mirowski et al., 2024; Ouyang et al., 2022; Askell et al., 2021; Bai et al., 2022; Sorensen et al., 2024). These approaches often lack the flexibility needed to adapt to the complexities of human-AI interactions, resulting in gaps in effectively aligning AI behavior with the varied expectations of everyday users (Klingefjord et al., 2024; McIntosh et al., 2024). As a result, Weidinger et al. propose to “account for relevant context” in the safety evaluation of LLMs to understand “who uses the AI systems, to what end and under which circumstances” (Weidinger et al., 2023b). As we discuss in § 5.4, when users try to re-align biased AI companions, they identify specific statements they deem harmful and respond to those statements based on the unique contexts in which they are situated, making the alignment process more relevant.
@@
Third, user-driven value alignment centers around personal and community expectations, values, and norms. Previous literature has highlighted how existing generalized value alignment processes—often termed global cultural value alignment—may overlook or even conflict with individual expectations and local community norms (Mirowski et al., 2024; Varshney, 2023). For example, the “helpful, honest, and harmless” framework (Askell et al., 2021) may directly conflict with the unique dynamics and expectations of the comedian community (Mirowski et al., 2024). Our research echoes these concerns, emphasizing the importance of tailoring the alignment process to better reflect the unique values and needs of users and their communities. Indeed, when some users detected biased statements, they deliberately chose not to correct the AI through technical means (e.g., modifying the settings) because doing so would disrupt their experiences with the AI companions (§ 5.4.1).
@@
Finally, user-driven value alignment supports iterative human-AI interactions. Unlike traditional user-driven algorithm auditing, user-driven value alignment involves users not only identifying AI biases but also actively participating in correcting and mitigating these harms. In the past, users struggled to directly intervene and modify system behavior (Shen et al., 2021; Li et al., 2023); however, our study shows that users can potentially influence the memory and future behavior of a specific AI companion through iterative interactions. This ease of engagement fosters the development of user-driven value alignment, where users play a crucial role in shaping AI systems and highlights the potential for building AI systems that better align with user values through ongoing processes.
6.2. Design Implications for Supporting User-Driven Value Alignment
@@
We discuss design implications for supporting more meaningful, effective, and safer user-driven value alignment. We also provide actionable design guidelines for practitioners to better support users in this process.
@@
Our research reveals that when users encounter biases and discrimination in AI companions, they often share the incidents via social media platforms. Some more experienced users also actively share strategies for addressing those harms and re-aligning their companions. Our research starts by collecting complaining posts users shared on a wide range of social media platforms. Through these activities, users—particularly those from marginalized social groups—engage in community-building efforts similar to what Nancy Fraser describes as “counterpublics” (Fraser, 2014; Shen et al., 2021). In those spaces, members of often marginalized social groups collectively participate in their own form of sensemaking, opinion formation, and consensus building. However, there is currently a lack of dedicated user communities focused on addressing biases in AI companions, which results in most users not having access to the strategies shared by more experienced users. Practitioners should consider tools and systems that support community collaboration. For example, dedicated spaces (e.g., a discussion forum) for users to discuss, share, and raise awareness about harmful AI behaviors. They may also consider augmenting the discussion platforms with features explicitly designed to support value alignment, such as an upvoting mechanism that allows users to highlight effective alignment strategies, collectively surfacing the most useful methods.
@@ 6.2.2. Enhancing effective user-driven value alignment via collaboration between experts and users
As we discussed in § 5.5, there remain gaps between users’ expectations and the perceived effectiveness of those alignment strategies, particularly for novice users. Towards this end, a combined approach that integrates expert-driven and user-driven value alignment should be considered. Expert-driven alignment provides a solid foundational framework, while user-driven alignment allows ongoing, context-sensitive adjustments. By merging these two approaches, AI systems can achieve a more dynamic and adaptable value alignment process. Empowering users through value-sensitive learning is key—AI should continuously learn and iterate based on user feedback, regularly updating personalized knowledge bases to better reflect diverse user values. When users flag harmful behaviors or provide improvement suggestions, this information should be integrated into the model’s optimization process. For practitioners, this could involve implementing bidirectional feedback mechanisms within the platform. On one side, an Expert-to-User channel allows experts such as developers to provide targeted guidance. For example, experts could suggest specific strategies or highlight areas where user input is particularly valuable. On the other side, a User-to-Expert channel empowers users to submit feedback using well-designed tools, such as structured forms or interactive interfaces that categorize and prioritize their suggestions. These mechanisms ensure that users receive meaningful support from experts while also providing experts with actionable insights to refine the system.
@@ 6.2.3. Facilitating safe user-driven value alignment via platform policies and affordances
First, for highly anthropomorphic AI agents like AI companions, platforms should have stronger constraints on their behavior to regulate what they should or should not express. For practitioners, this might be achieved by improving the instruction-following capabilities of the underlying LLM (Guo et al., 2024; Mu et al., 2023), optimizing output filtering mechanisms (Glukhov et al., 2023), or establishing a more comprehensive user-specific interaction database to store and reference positive and negative interaction examples.
Second, platforms should be designed to ensure that users are fully aware of the resources available in case of a harmful value misalignment. It’s crucial that users know they have the option to re-align their AI companions. This approach goes beyond simply providing re-alignment tools; it emphasizes the importance of clearly communicating these options to users, empowering them to take corrective action rather than leaving the platform out of frustration. Practitioners should consider designing user-friendly interfaces that highlight these resources and provide step-by-step guidance on how to use re-alignment tools effectively. Additionally, incorporating proactive notifications or prompts to inform users about these options can enhance accessibility and encourage engagement.
Third, platforms should offer robust support when users need to re-align their AI companions, rather than leaving them to rely solely on intuition—especially since some strategies may be less effective, particularly for novice users. This support could include value alignment assistants, platforms for sharing effective strategies, and other resources. These tools should be tailored to the user’s conceptual understanding of the AI companion. For example, discussing technical details like training data might not resonate with users who view their AI companion as a “baby” and could disrupt their experience, as mentioned in § 5.3.2. Additionally, it’s essential that platforms clearly communicate the steps, tools, and resources available for re-alignment, ensuring that users understand how to restore their relationship with the AI after a misalignment rather than abandoning the application entirely.
Fourth, when designing AI systems, it is essential to consider both users’ immediate needs and how to establish and maintain long-term emotional connections and behavior improvement mechanisms. Practitioners should provide more effective and low-risk solutions, focusing on users’ mental health and avoiding intense emotional confrontation.
6.3. Concerns and Challenges of User-Driven Value Alignment
@@ 6.3.1. Concerns about malicious users
Malicious users may pose risks to AI systems, a concern that has been a focus in the field of responsible AI. A typical example is Microsoft’s chatbot Tay, which quickly exhibited racist behavior after interacting with users on Twitter (Schwartz, 2019). While our research underscores the many advantages of enabling user-driven value alignment, it is crucial to acknowledge the associated risks and challenges. As discussed in Sections 5.4, some users may maliciously exploit the AI’s Chameleon Effect (Chartrand and Bargh, 1999), where the AI mimics human behavior to manipulate and contaminate the training data. They might use insulting or threatening language, causing the AI to replicate similar harmful behavior in future interactions. Indeed, one key advantage of expert-driven, top-down value alignment is its ability to prevent harm through strict oversight, with platforms being held accountable for ensuring that AI systems operate safely and ethically. This approach centralizes responsibility, allowing for more consistent and controlled implementation of safeguards. Shifting this responsibility to users and smaller user communities can inevitably introduce potential risks. To mitigate these risks, it is essential to explore mechanisms that combine both expert-driven and user-driven value alignment (§ 6.2.2) and to foster greater community accountability (§ 6.2.1).
@@
6.3.2. Concerns about potential harms and ethical implications of involving users in correcting AI biases
There are also concerns about the potential harms and ethical implications of placing the responsibility of correcting AI biases on users. On the one hand, shifting the responsibility of addressing AI bias onto users may allow platforms and developers to evade their obligations to reduce bias and prevent discriminatory outputs (Eubanks, 2018). This approach risks exacerbating inequalities, as users with greater technical knowledge or emotional resources are more likely to participate, while those with fewer resources may become further marginalized. Therefore, user involvement must be voluntary and prioritize users’ well-being rather than exploiting their efforts to compensate for flaws in AI system design (Morreale et al., 2023).
On the other hand, we also need to pay attention to the psychological and emotional burden on users when directly engaging in mitigating harmful AI systems. Research on content moderation shows that moderators often face psychological harm and emotional labor when dealing with harmful content (Dosono and Semaan, 2019; Steiger et al., 2021; Wohn, 2019). Our study found that with the advancement of LLMs, interactions with highly anthropomorphized AI agents (e.g., AI companions) may further intensify these negative effects (§ 5.1). For instance, prolonged interaction with biased AI systems may desensitize users to harmful outputs, potentially normalizing biases (Kingsley et al., 2024). The emotional reliance of users on AI companions raises another layer of concern. Overdependence on AI for emotional support may lead to social withdrawal and feelings of isolation, particularly for users already experiencing loneliness, sadness, or social marginalization. These users might unintentionally prioritize developing relationships with AI over real-world social interactions, further weakening their societal engagement (Smith and Kelly, 2024). To help alleviate those harmful impacts, practitioners should proactively draw on existing methods such as combining workplace strategies, clinical practices, and technological interventions to create comprehensive well-being support systems (Steiger et al., 2021; Zhang et al., 2024; Li and Zhang, 2024).
6.3.3. When the design implication is not to re-align (Baumer and Silberman, 2011)?
@@
There are questions about whether we should choose to re-align certain harmful AI systems at all. User-driven value alignment can be considered a type of “repair” work (Velkova and Kaun, 2021) that users perform in their daily interactions with AI companions to reduce biases. However, researchers have also pointed out that in some cases, “refusal”—rather than “repair”—should be considered a more appropriate response to mitigate the potential harm that systems can inflict on users (Ganesh and Moss, 2022; Zong and Matias, 2024). For example, although users generally enjoy interacting with AI companions, the high degree of anthropomorphism may lead to over-reliance on these systems for emotional support, neglecting real-life interpersonal relationships (Zimmerman, 2009). This phenomenon is particularly concerning for vulnerable or marginalized users. In such cases, reducing users’ dependence on AI companions, rather than focusing solely on re-aligning these systems to mitigate biases, can be a more effective way to prevent harm. Additionally, some AI systems require significant and sustained user intervention to address issues like bias, which can be resource-intensive and unsustainable. When the effort and resources required for re-alignment significantly outweigh the potential benefits, limiting or discontinuing such systems may be a more practical and beneficial option.
@@ 6.4. Limitations and Future Work
As the first work conceptualizing and studying user-driven value alignment, this research is highly exploratory, and there are a few important limitations. First, our research relies on self-reported data from interviews and social media posts, which biases may influence. The study primarily focuses on users’ perspectives, so insights from other stakeholders involved in the value alignment process would be valuable for gaining a more holistic understanding of the phenomenon. Second, the qualitative nature of the study and its limited sample size limit our ability to generalize the findings broadly. While this research outlines seven alignment strategies and probes around the perceived effectiveness of those strategies based on qualitative interview data from users’ perspectives, their actual effectiveness and long-term impact have yet to be thoroughly evaluated in complicated, real-world contexts. Future research could explore which strategies are more effective under specific conditions via complementary methods, such as controlled experiments or longitudinal studies. Third, this study focuses on AI companions that may not represent all AI systems. It focuses on biases and discriminatory AI behaviors, which may not encompass all the problematic behaviors that a value alignment process should address. It will be valuable to explore whether and how useful the concept of user-driven value alignment is in terms of studying other types of value alignments in different AI systems. Fourth, this study emphasizes biases and harmful behaviors in LLMs; future research should consider multimodal risks and improve mechanisms for detecting and correcting multimodal biases.